# How quantum and evolutionary algorithms can help each other: two examples

### v1

@@ 2.2 Designing entangling quantum circuits
In this section, we present the results obtained with an evolutionary algorithm as a tool to create a circuit that entangles qubits. The strength of the entanglement is quantified with the Mayer-Wallach measure which plays the role of a fitness function. We consider systems of 3, 4, and 5 qubits, and the performance of the algorithm is explored by varying mutation rates and numbers of gates. We ran experiments for 500 generations, repeating it 50 times. The outcomes are shown as the average fitness along with the standard error, and the best fitness achieved, also with its standard error.
Figure 5: Top: Evolutionary optimization of three-qubit quantum circuits with three gates using the Meyer-Wallach entanglement measure [21] as the fitness function. (left) Plots show mean and best fitness across generations for different mutation rates: 5%, including a third-order polynomial fit to the mean fitness. (right) Best fitness comparison with varying mutation rates and gate numbers. Bottom: (left) Fitness outcomes for different gate numbers at a constant 10% mutation probability, averaged over 50 runs with standard error bars. (right) An example circuit achieves a high fitness score of 0.999 with only three gates and 10% mutation probability.
@@ 2.2.1 Three-qubit circuits
Since the best fitness scores for the quantum circuits approached unity, evolutionary algorithms indeed serve as a toolbox of methods to design highly entangling three-qubit circuits. For a three-qubit system (three lists of integer gates for each qubit), a mutation probability of 1/9 (11.11%) is needed to replace at least one individual population consisting of nine integer lists. Therefore, a 10% mutation rate, which can at least replace one individual population in each generation can be considered an optimal probability for the mutation rate. Experimentally, the optimal mutation rate was found to be 10%, which struck a balance between exploring different solutions in the search space and exploiting the best solutions as shown in Figure 5 (Top: right). In addition, a lower mutation rate is capable of striking a balance between exploration and exploitation, leading to a higher average fitness score across all runs. However, the mutation rate of 10% is more effective in building and generating the circuit that performs best in each run due to the higher degree of exploration it allows for the entanglement.
We also varied the of gates in the quantum circuit in the range from 3 to 12 keeping the optimal mutation rate at 10%. The result is the fitness score decreased as the number of gates in the circuit increased as shown in Figure 5 (Bottom: Left). These results suggest that as the quantum circuit becomes more complex with an increasing number of gates, the entanglement capability of the circuit is reduced, indicating that the circuit is less efficient at performing the desired task. This suggests that there may be a trade-off between circuit complexity and performance. Therefore choosing the optimal number of gates can provide the best balance between the complexity of the quantum circuit and the performance.
The best three-qubit entangling circuit with three gates is shown in Figure  5 (Bottom: right). It corresponds to the fitness score of 0.999. Similarly, a three-qubit quantum circuit with 12 gates also achieved a fitness value of 0.999. In Supplementary Note II, the detailed analyses of the state vector and reduced density matrices from optimal circuits in a 500-generation are presented. Table 1 in Supplementary Note shows these details for the state vector probability of the respective density matrices. The methodologies used for calculating the probabilities of quantum states and the reduced density matrices for the three-qubit circuits, along with an assessment of their entanglement using the Mayer-Wallach entanglement measure, are given in Supplementary Note III.
@@ 2.2.2 Four and five-qubit circuits
Figure 6: Top: (left) Four-qubit quantum circuits with four gates and a 10% mutation probability. The mean fitness (green line) and its shaded standard error, along with the mean of the best fitness (red line) and its shaded standard error, plotted against the number of generations. (right) Comparison of the best fitness generated for five different numbers of gate sets in a four-qubit circuit using a 10% mutation rate and 20 chromosomes. Bottom: Evolutionary generation of four-gate four-qubit circuits with MW-entanglement fitness scores using a 10% mutation probability with a fitness score of 0.9996.
@@
In this section, we extended the designing of quantum circuits to advance our findings to include both four and five-qubit systems. The findings are shown in Figures 6 and 7 for four and five qubit quantum circuits, respectively. Notably, Figures (Top: Left) 6 and 7 show the mean fitness (illustrated by a green line) alongside its standard error (denoted by a shaded area), in addition to the average of the best fitness (represented by a red line) and its standard error (also shaded). A third-order polynomial fit applied to the mean fitness is shown by a blue dashed line. These figures representations demonstrate that the average fitness score approximates 0.6, while the mean of the highest fitness score approaches 0.8.
@@
Further experiment with an optimal mutation probability of 10% and 5% for four and five qubit circuits, respectively, was conducted focusing on circuits with varying numbers of gates as shown in Figures (Top: right) 6 and 7. Through an analysis of mutation rates of 3%, 5%, and 15%, we found that increasing the mutation rate beyond 5% results in a decrease in the average fitness value. This indicates that excessive randomness in the evolutionary process can have a negative impact on circuit performance. However, it was also observed that the average of the best fitness value was not affected by higher mutation rates. This highlights the importance of selecting an appropriate mutation rate for obtaining the quantum circuits of maximal entanglement.
The visualization of the quantum circuit with four and five qubits incorporating four and five gates, respectively, achieving a fitness score of 0.999, is shown in Figures (Bottom) 6 and 7. Additionally, the five-qubit quantum circuit with 12 gates yielded a fitness score of 0.199, signifying a decrement in entanglement capabilities with an increase in the number of gates. Detailed outcomes for the most entangled four and five-qubit quantum circuits, including the state vector and reduced density matrices, are elaborated in Section II of the Supplementary Note. These results pertain to the optimal five-qubit circuit identified across 500 generations, encompassing circuits with both five and twelve gates, as illustrated in Tables 2 and 3 (Section II of the Supplementary Note).
@@
Figure 7: Evolutionary optimization of five-qubit quantum circuits: Top: (left) Five gates with 3% mutation, (right) Five gates with 5% mutation, and (Bottom) Varying gate numbers with 5% mutation. The plots showcase mean and best fitness scores (green and red lines, respectively) with standard error against the number of generations. The blue dashed lines represent third-order polynomial fits to the mean fitness. Bottom: fitness scores for circuits with five gates (score: 0.999).
@@ 3 Discussion and conclusions
In this study, we explore the two examples of evolutionary algorithms in two distinct yet interconnected areas of quantum computing. Initially, we introduce a framework for generating quantum circuits that replicate specific cellular automata rules using genetic algorithms [18]. The implementation code is available at GitHub repository https://github.com/Overskott/Quevo. This framework effectively evolves quantum circuits for various CA rules, from deterministic to stochastic types. We improved our original framework by increasing the number of generations and runs, enhancing performance and more consistent fitness function outcomes. This initial investigation set the stage for more nuanced applications of evolutionary algorithms in quantum circuit design, suggesting future research paths.
To conclude, building upon this groundwork, we have demonstrated the effectiveness of evolutionary algorithms in generating highly entangled quantum circuits, emphasizing their role in calculating reduced density matrices for three-, four-, and five-qubit systems. These circuits displayed a significant degree of entanglement, measured by the Mayer-Wallach entanglement measure. By tracing out all other qubits, the study illustrated how reduced density matrices could reveal detailed information about the circuit’s entanglement. In this context, the algorithm’s role was to optimize quantum gates and their connectivity, leading to the creation of quantum circuits.
This contribution is, as far as we have checked, novel in its approach to optimizing quantum circuits for entanglement. Moreover, the simulator provided by Qiskit enabled the evaluation of these circuits, gates, and connections, showcasing the potential of this approach in solving complex entanglement optimization problems. Furthermore, we highlighted the importance of circuit depth in determining the entanglement capability of a quantum circuit, and we found that increasing the circuit depth decreases its entanglement capability.
@@
Another important conclusion is that, an optimal mutation rate was identified for quantum circuits with different qubits and gates, marking a significant stride in optimizing quantum circuits for entanglement generation. To note that the KL-fitness score measures the discrepancy between the distributions of the output states and the target states. In our case, these distributions are discrete, each corresponding to the eight different initial states of the three qubits. If the distributions completely match the fitness function, the KL-value is zero. A decreasing fitness value close to zero indicates an increase in the efficiency/entanglement of the circuits. Our findings show that with mutation rates up to 20%, the optimization process effectively explores the solution space, leading to better circuit designs with lower fitness scores. However, mutation rates beyond 20% introduce excessive randomness, which disrupts the optimization process and results in higher fitness scores, indicating less efficient/entangled circuits. All mutation rates result in a decrease in fitness scores over generations, indicating improvement. However, mutation rates of 5%, 10%, and 20% stabilize more consistently compared to higher rates. The right panel box plots show that while higher mutation rates (30%, 50%) lead to decreased fitness scores, they introduce greater variability and less stability. These results highlight that while higher mutation rates can still improve circuit efficiency, the optimal balance between exploration and stability during the "evolution" of the algorithm is achieved at a 20% mutation rate.
@@
Finally, some last points for discussion. This paper opens avenues for future investigations into advanced evolutionary algorithms and alternative fitness functions, such as Von Neumann entropy [22] and Schmidt measures [23, 24], to maximize entanglement in more complex quantum circuits. Future work could explore the scalability of these methods to larger quantum circuits, composed by more than five qubits. The fitness function here introduced, measuring the entanglement between qubits could be extended to multi-dimensional optimization scheme, incorporating also measures of error and noise in quantum circuits. Moreover, it would also be of interest to adapt the method here proposed and integrate it in hybrid optimization approaches. Future research could explore the maximization of entanglement in quantum circuits, drawing inspiration from recent advancements in benchmarking highly entangled states on a 60-atom analog quantum simulator [25] and leveraging non-locality as a resource in quantum computing demonstrated by loophole-free Bell inequality violations with superconducting circuits [25]. All in all, applying optimized quantum circuits to real-world problems in quantum cryptography [26, 27], quantum simulation [27, 28], and quantum machine learning [29], as well as testing them on actual quantum hardware, could further demonstrate their practical utility and effectiveness [28].
4 Methods
Our study implements two distinct methods for designing quantum circuits. The first method utilizes cellular automata rules combined with the KL fitness function. Here we advanced the original framework from Ref. [18] for generating quantum circuits that replicate quantum circuits with evolutionary algorithms. Upon this groundwork the second method focuses on designing quantum circuits for maximum entanglement, using the Mayer-Wallach measure and von Neumann entropy for entanglement evaluation. The comprehensive details regarding the implementation of evolutionary quantum algorithms for both methodologies are thoroughly delineated in Supplementary Note I.
4.1 Evolutionary algorithms: Mutations and fitness function
The evolutionary algorithm initiates with a specific assigned number of chromosomes, each populated with a set number of quantum gates within the circuit. This initial population is then subject to evolution, where the fittest chromosomes are identified, iteratively refining the existing ones, and introducing the best-fit chromosome at the forefront of each new generation. This evolutionary cycle continued for a predetermined number of generations, with each generation’s best chromosome becoming the parent for the next. Throughout this process, the chromosomes are subjected to mutations to enhance the optimization. The mutation process is introduced in two different forms: either by swapping existing gates in the chromosome with randomly selected ones from the gate pool, or by entirely replacing the chromosome to identify the best four parental candidates suitable for the next generation. The mutation process is controlled by a set of probabilities. The four best chromosomes are preserved as "elites" without changes, whereas the remaining chromosomes are selected for evolution based on their fitness, with selection changes directly linked to their performance.
To evaluate the performance of the quantum circuits, we measure only the first qubit (q0) and consider the fitness function as the Kullback-Leibler (KL) divergence, which measures the difference between two probability distributions and has been used in other works [30, 31] as a fitness function:
4.2 Assessing entanglement in basic quantum states of qubits
Consider, for the start, the state of a system of two-qubit configurations of a quantum system. The most general form of a state for such a system can be written as:
$\begin{split}|\psi\rangle=\alpha_{00}\ket{00}+\alpha_{01}\ket{01}+\alpha_{10}\ket{10}+\alpha_{11}\ket{11}\,\,,\end{split}$ (2)
$|\psi\rangle=\frac{1}{\sqrt{2}}\left({\ket{00}+\ket{11}}\right)\,.$ (4)
[…]
In this state, when measuring one qubit, there is no certainty about the state of the other qubit. Namely, there is a probability of $1/2$ that the other qubit is in state 00 and probability $1/2$ that it is in state $1$.
In general, suppose we have two quantum systems $Q_{1}$ and $Q_{2}$. These systems are said to be entangled if the values of certain properties of system $Q_{1}$ (accessed with measurement) are correlated with the values that those properties will assume for system $Q_{2}$. The simplest forms of entangled state are known as the Bell states [33], which enable tasks such as quantum cryptography[34], super-dense coding[35], teleportation[36], and entanglement swapping [37].
For example, the measure of expressibility can be calculated by taking the Kullback-Leibner divergence between the estimated fidelity distribution and that of the Haar-distributed ensemble [39]. In this paper, we considered two measures, the Meyer-Wallach (MW)[40] and the entanglement entropy (EE)[41, 42, 43]. The Meyer-Wallach entanglement measure is popular because of its scalability and ease of computation. It considers a pure state of $n$ qubits if the form [21]:
$Q(\ket{\psi})=\frac{4}{n}\sum_{j=1}^{n}D(\ket{\hat{u}^{k}},\ket{\hat{v}^{k}})\,,$ (6)
where |$\ket{\hat{u}^{k}}$ and |$\ket{\hat{v}^{k}}$ denote non-normalized vectors belonging to the complex vector space $\mathbf{C^{2n-2}}$, where $n$ denoting the number of qubits in the quantum system. These vectors are obtained by projecting the pure state |$\ket{\psi}$ of the $n$-qubit system onto the local basis states of the $k^{th}$ qubit, where $k\in{1,2,...,n}$. The symbol "$\;\hat{}\;$" above the vectors signifies their non-normalized status. The function $D(\ket{\hat{u}^{k}},\ket{\hat{v}^{k}})$ measures a distance between the two vectors |$\ket{\hat{u}^{k}}$ and |$\ket{\hat{v}^{k}}$, measured through the generalized cross-product:
$D(\ket{\hat{u}^{k}},\ket{\hat{v}^{k}})=\sum_{i<j}|{\hat{u}_{i}^{k}\hat{v}_{j}^{k}-\hat{u}_{j}^{k}\hat{v}_{i}^{k}}|^{2}\,.$ (7)
Since the purity of the state of qubit $k$ is given by $Tr[\rho_{k}^{2}]=\braket{\hat{x}^{k}}{\hat{x}^{k}}^{2}+\braket{\hat{y}^{k}}{\hat{y}^{k}}^{2}$, we obtain [21]
$D(\ket{\hat{x}^{k}},\ket{\hat{y}^{k}})=\sum_{i<j}|{\hat{u}_{i}^{k}\hat{v}_{j}^{k}-\hat{u}_{j}^{k}\hat{v}_{i}^{k}}|^{2}$ (8)
[…]
$Q(\ket{\psi})=2\left(1-\frac{1}{n}\sum_{k=0}^{n-1}Tr[\rho_{k}^{2}]\right),$ (9)
which is the usual definition of the Meyer-Wallach measure. A detailed derivation and discussion of this measure, especially its implications for both pure and mixed states, are provided in Supplementary Note IV.
$S(\rho_{A})=-\mathrm{Tr}(\rho_{A}\log_{2}\rho_{A})\,,$ (10)
In what concerns implementation, we have used the mathematical formulations presented in Equations (9) and (10) in a framework previously developed (https://github.com/Overskott/Quevo). It is further aligned with the methodologies outlined in Ref.[18]. The evolutionary algorithms (EAs) have emerged as a promising tool for quantum circuit design [45, 6, 7]. So our implementation of EAs for quantum circuit design aims to maximize entanglement using two specified methods. The framework employs an evolutionary algorithm to automatically generate quantum circuits that satisfy defined properties specified through a fitness function. The evolution properties can be controlled by tuning parameters such as initial population, number of generations, probability of mutation operator, number of gates in the circuit, and the chosen fitness function. Details of the implementation process are available in Supplementary Note III.
5 Data availability
@@
The datasets used and/or analyzed during the current study are available in the QUEVO1 repository, https://github.com/shailendrabhandari/QUEVO1.

### v2

@@ 3.2 Designing entangling quantum circuits
In this section, we employ an evolutionary algorithm to design quantum circuits that generate entangled states. To establish a proof of concept, we will specifically address systems of three, four, and five qubits, using the MW entanglement measure as a fitness function. We analyze these qubit numbers separately to explore how entanglement generation varies with system size, a critical factor in quantum circuit design. In the end, we discuss how the feasibility of our framework scales with the complexity of larger quantum circuits. Our experiments, conducted over 500 generations with 50 runs each, reveal how these factors influence entanglement optimization across different system sizes.
a)Refer to captionb)
Refer to captionc)Refer to captiond)
Figure 5: a) Evolutionary optimization of three-qubit quantum circuits with three gates using the MW entanglement measure [15] as the fitness function. b) Plots show mean and best fitness across generations for different mutation rates: 5%, including a third-order polynomial fit to the mean fitness. c) Best fitness comparison with varying mutation rates and gate numbers. d) Fitness outcomes for different gate numbers at a constant 10% mutation probability, averaged over 50 runs with standard error bars. An example circuit achieves a high fitness score of 0.999 with only three gates and 10% mutation probability.
@@ 3.2.1 Three-qubit circuits
Since the best fitness scores for the quantum circuits approached unity, EAs indeed serve as a toolbox of methods to design highly entangling three-qubit circuits. For a three-qubit system (three lists of integer gates for each qubit), a mutation probability of 1/9 (11.11%) is needed to replace at least one individual population consisting of nine integer lists. Therefore, a 10% mutation rate, which can at least replace one individual population in each generation, can be considered an optimal probability for the mutation rate. Experimentally, the optimal mutation rate was found to be 10%, which struck a balance between exploring different solutions in the search space and exploiting the best solutions as shown in Figure 5 b). In addition, a lower mutation rate is capable of striking a balance between exploration and exploitation, leading to a higher average fitness score across all runs. However, the mutation rate of 10% is more effective in building and generating the circuit that performs best in each run due to the higher degree of exploration it allows for the entanglement.
We also varied the number of gates in the quantum circuit in the range from 3 to 12, keeping the optimal mutation rate at 10%. The result is that the fitness score decreased as the number of gates in the circuit increased, as shown in Figure 5 c). These results suggest that as the quantum circuit becomes more complex with an increasing number of gates, the entanglement capability of the circuit is reduced, indicating that the circuit is less efficient at performing the desired task. This suggests that there may be a trade-off between circuit complexity and performance. Therefore, choosing the optimal number of gates can provide the best balance between the complexity of the quantum circuit and the performance.
The best three-qubit entangling circuit with three gates is shown in Figure 5 d). It corresponds to the fitness score of 0.999. Similarly, a three-qubit quantum circuit with 12 gates also achieved a fitness value of 0.999. In Supplementary Note III, the detailed analyses of the state vector and reduced density matrices from optimal circuits in a 500-generation circuit are presented. Table 2 in the Supplementary Note shows these details for the state vector probability of the respective density matrices. The methodologies used for calculating the probabilities of quantum states and the reduced density matrices for the three-qubit circuits, along with an assessment of their entanglement using the MW entanglement measure, are given in Supplementary Note III.
a)Refer to captionb)
Refer to captionc)
Figure 6: a) Four-qubit quantum circuits with four gates and a 10% mutation probability. The mean fitness (green line) and its shaded standard error, along with the mean of the best fitness (red line) and its shaded standard error, are plotted against the number of generations. b) Comparison of the best fitness generated for five different numbers of gate sets in a four-qubit circuit using a 10% mutation rate and 20 chromosomes. c) Evolutionary generation of four-gate four-qubit circuits with MW-entanglement fitness scores using a 10% mutation probability with a fitness score of 0.9996.
@@ 3.2.2 Four and five-qubit circuits
In this section, we extended the design of quantum circuits to advance our findings to include both four and five-qubit systems. The findings are shown in Figures 6 and 7 for four and five-qubit quantum circuits, respectively. Notably, Figures 6 a) and 7 a) show the mean fitness (illustrated by a green line) alongside its standard error (denoted by a shaded area), in addition to the average of the best fitness (represented by a red line) and its standard error (also shaded). A third-order polynomial fit applied to the mean fitness is shown by a blue dashed line. These figure representations demonstrate that the average fitness score approximates 0.6, while the mean of the highest fitness score approaches 0.8.
@@
Further experiment with an optimal mutation probability of 10% and 5% for four and five qubit circuits, respectively, was conducted, focusing on circuits with varying numbers of gates as shown in Figures 6 b) and 7 b). Through an analysis of mutation rates of 3%, 5%, and 15%, we found that increasing the mutation rate beyond 5% results in a decrease in the average fitness value. This indicates that excessive randomness in the evolutionary process can have a negative impact on circuit performance. However, it was also observed that the average of the best fitness value was not affected by higher mutation rates. This highlights the importance of selecting an appropriate mutation rate for obtaining the quantum circuits of maximal entanglement.
a)Refer to captionb)
Refer to captionc)
@@
Figure 7: Evolutionary optimization of five-qubit quantum circuits: a) Optimization of five-qubit circuits with five gates and a 5% mutation rate. The plots show the mean (green line) and best (red line) fitness scores, with standard error bars plotted against the number of generations. The blue dashed lines represent third-order polynomial fits to the mean fitness scores. b) Evolutionary optimization with varying gate numbers, all with a 5% mutation rate. c) Fitness scores for circuits with five gates, achieving a score of 0.999.
The visualization of the quantum circuit with four and five qubits, incorporating four and five gates, respectively, achieving a fitness score of 0.999, is shown in Figures 6 c) and 7 c). Additionally, the five-qubit quantum circuit with 12 gates yielded a fitness score of 0.199, signifying a decrement in entanglement capabilities with an increase in the number of gates. Detailed outcomes for the most entangled four and five-qubit quantum circuits, including the state vector and reduced density matrices, are elaborated in Section II of the Supplementary Note. These results pertain to the optimal five-qubit circuit identified across 500 generations, encompassing circuits with both five and twelve gates, as illustrated in Tables 3 and 4 (Supplementary Note IV).
@@ 4 Discussion and conclusions
In this study, we explore the two examples of EAs in two distinct yet interconnected areas of quantum computing. Initially, we introduce a framework for generating quantum circuits that replicate specific cellular automata rules using genetic algorithms [13]. The implementation code is available at the GitHub repository https://github.com/Overskott/Quevo. This framework effectively evolves quantum circuits for various CA rules, from deterministic to stochastic types. We improved our original framework by increasing the number of generations and runs, enhancing performance, and achieving more consistent fitness function outcomes. This initial investigation set the stage for more nuanced applications of EAs in quantum circuit design, suggesting future research paths.
Building upon this groundwork, we have demonstrated the effectiveness of EAs in generating highly entangled quantum circuits, emphasizing their role in calculating reduced density matrices for three-, four-, and five-qubit systems. These circuits exhibited significant entanglement, as quantified by the MW entanglement measure. At the same time, as the number of qubits increases, so does the complexity of entanglement—from well-known three-qubit states like GHZ or W states [55] to more intricate multipartite structures in four- and five-qubit systems [37]. This progression allows us to assess the scalability of our evolutionary approach, identify optimal circuit configurations, and evaluate the trade-off between circuit complexity (e.g., gate count) and entanglement performance. Moreover, they are of practical relevance, since studying three to five qubits aligns with the capabilities of near-term quantum hardware, where noise and decoherence limit the practical utility of large qubit counts.
However, while the sizes chosen in our work are sufficient for proof-of-concept demonstrations, our findings can now be extended toward complexity levels beyond the sizes addressed in this work. Although with limited access, current quantum hardware, such as IBM’s devices with several hundred qubits [1], highlights the need to assess how our methods perform beyond this range. Scaling to larger systems, with 10–20 qubits or more, exponentially increases the search space for evolutionary optimization, which is already slow and computationally costly by design. This scalability challenge, when combining quantum and EAs, is compounded by the trade-off between circuit complexity and entanglement performance. In this respect, our results show that entanglement capability decreases with increasing circuit depth in small systems. Specifically, a five-qubit circuit with 12 gates scores 0.199 versus 0.999 with five gates. In larger systems, this trade-off is likely to intensify, as deeper circuits may be required to achieve entanglement across more qubits, yet excessive depth could further degrade performance due to accumulated errors and decoherence.
To stress that, our focus on small system sizes is motivated by the constraints of near-term quantum hardware, where noise and decoherence limit the practical use of large qubit counts, rendering optimized small circuits highly relevant for current applications. Nevertheless, we acknowledge that this restricts the immediate applicability of our findings to larger-scale quantum computing scenarios. To address this, future work could leverage hybrid classical-quantum optimization techniques or execute the EAs directly on quantum hardware, potentially offering speed advantages and enabling scalability to larger systems. Such advancements could extend our approach to larger, more realistic quantum systems, enhancing their practical applicability in settings where hundreds of qubits are available.
@@
Another key finding is the identification of an optimal mutation rate for quantum circuits across different qubit sizes and gate counts, marking a significant step in optimizing entanglement generation. To note that the $D_{KL}$ fitness score measures the discrepancy between the distributions of the output states and the target states. In our case, these distributions are discrete, each corresponding to the eight different initial states of the three qubits. If the distributions completely match the fitness function, the $D_{KL}$ value is zero. A decreasing fitness value close to zero indicates an increase in the efficiency/entanglement of the circuits. Our findings show that with mutation rates up to 20%, the optimization process effectively explores the solution space, leading to better circuit designs with lower fitness scores. However, mutation rates beyond 20% introduce excessive randomness, which disrupts the optimization process and results in higher fitness scores, indicating less efficient/entangled circuits. All mutation rates result in a decrease in fitness scores over generations, indicating improvement. However, mutation rates of 5%, 10%, and 20% stabilize more consistently compared to higher rates. The right panel box plots show that while higher mutation rates (30%, 50%) lead to decreased fitness scores, they introduce greater variability and less stability. These results highlight that while higher mutation rates can still improve circuit efficiency, the optimal balance between exploration and stability during the "evolution" of the algorithm is achieved at a 20% mutation rate.
@@
Finally, some last points for discussion. This paper opens avenues for future investigations into advanced EAs and alternative fitness functions, such as Von Neumann entropy [35] and Schmidt measures [21, 20], to maximize entanglement in more complex quantum circuits. Future work could explore the scalability of these methods to larger quantum circuits, composed of more than five qubits. The fitness function here introduced, measuring the entanglement between qubits, could be extended to a multi-dimensional optimization scheme, incorporating also measures of error and noise in quantum circuits. Moreover, it would also be of interest to adapt the method proposed here and integrate it into hybrid optimization approaches. Future research could explore the maximization of entanglement in quantum circuits, drawing inspiration from recent advancements in benchmarking highly entangled states on a 60-atom analog quantum simulator[49] and leveraging non-locality as a resource in quantum computing demonstrated by loophole-free Bell inequality violations with superconducting circuits[46]. Ultimately, applying our optimized quantum circuits to real-world problems, such as quantum cryptography[39, 43], quantum simulation[43, 46], and quantum machine learning[44]—could further demonstrate their practical utility and effectiveness. In gate-model quantum computing frameworks, such as scalable distributed systems[26] or circuit depth reduction efforts[25], our EAs could enhance efficiency by minimizing gate counts while maximizing entanglement, aligning with near-term hardware needs. Similarly, in a quantum internet context[27], our highly entangled circuits could support protocols like quantum key distribution or teleportation, contributing to distributed quantum networks. Also, one could explore the scalability of our approach to larger quantum circuits, building on the experimental advancements demonstrated by Chen et al. [18] and the theoretical frameworks proposed by Arrighi et al. [5]. Additionally, integrating noise resilience techniques into our evolutionary algorithm framework could further enhance the stability and performance of quantum circuits in noisy environments.
Data and code availability
@@
The datasets and the implementation code used and/or analyzed during the current study are available in the QUEVO1 repository, https://github.com/shailendrabhandari/QUEVO1.